OpenStudy (anonymous):

What is the probability of picking 5 puppies of which 3 are ale from a group of 5 male and 4 female

OpenStudy (zarkon):

10/21

OpenStudy (anonymous):

could you explain how you got htat

OpenStudy (zarkon):

hypergeometric distribution

OpenStudy (anonymous):

The total group size is (5+4) = 9 We need to select 5 from this group of 9 The ways to do this are 9C5 = 126

OpenStudy (anonymous):

Now, we are interested in the case where there are 3males (and thus by duduction 2 females). So we need to choose 3 males out of the 5males available. We also need to select 2 females out of the 4 females available The ways we can satisfy the condition is thus ( 5C3 x 4C2) = 10x6=60 Thus the answer is 60/126 = 10/21